OpenStudy (anonymous):

There are 7 books on the floor. The number of books on the shelf is 2 more than 3 times the number of books on the floor. How many books are on the shelf?

OpenStudy (idealist10):

3*7+2=21+2=23 23 books are on the shelf.

OpenStudy (anonymous):

thanks

OpenStudy (idealist10):

you're welcome!